OpenStudy (anonymous):

Simplify the expression using the properties of exponents: (3y)^-3

OpenStudy (anonymous):

\[\frac{1}{(3y)^3}=\frac{1}{27y^3}\]\]

OpenStudy (anonymous):

1/(27 y^3)

OpenStudy (anonymous):

now ask your math teacher why \[\frac{1}{27y^3}\] is "simpler" than \[(3y)^{-3}\]

OpenStudy (anonymous):

because frankly there is nothing "simpler" about either one. one is written in parentheses on one line with a negative exponent. the other is written as a fraction with positive exponents. they each have exactly 6 symbols by my count.